News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Sunday 12 September 1999 12.00pm EST

During the last 24 hours the on-board load (P13) began operation and a number of ACIS tests were conducted in parallel. The ACIS test to lower the focal plane temperature to -110 C resulted in a significant improvement of the front-side illuminated chip energy resolution. A further test at -120 C is currently in progress. The SIM was positioned with ACIS at the focal plane for the current radiation passage in order to test the theory that resolution degradation is related to SIM position during radiation passage.

An anomaly related to the LETG insertion commanded by the load was detected during the communication pass at 7am 9/12. The grating insertion near the start of the load did not occur and the flight software aborted the insertion sequence. The load continued to run as expected. Analysis of the dump data indicated that the Mechanism Control Electronics (MCE) was not turned off at the completion of the previous retraction and is thought to be the cause. The behavior of conditions. Current plan is to perform the LETG insertion via real-time commanding and continue analysis with focus on the software timing.

Observations of Capella are planned with HRC-S and LETG in conjunction with the Hubble Space Telescope, EUVE, BeppoSAX and the Very Large Array (VLA) over the next 24 hours.





Return to Status Reports